The probability function of a discrete random variable X is as follows:

Values of X: x -4 -2 0 2 4

P(x) k 2k 3k 5k 6k


i) Find the value of k.

ii) Find the probability of the value of X exactly one.

iii) Find the probability of the value of X between -2 and 2.

iv) Estimate expected value and standard deviation of X.



1
Expert's answer
2022-02-01T16:51:41-0500

i)


"\sum_ip_i=1"

"k +2k+ 3k+ 5k+ 6k=1"

"k=1\/17"

ii)


"P(X=1)=0"

iii)


"P(-2\leq X\leq2)=P(X=-2)+P(X=0)"

"+P(X=2)=2\/17+3\/17+5\/17=10\/17"

"P(-2<X<2)=P(X=0)=3\/17"

iv)


"E(X)=\dfrac{1}{17}(-4)+\dfrac{2}{17}(-2)+\dfrac{3}{17}(0)+\dfrac{5}{17}(2)"

"+\dfrac{6}{17}(4)=\dfrac{26}{17}"


"E(X^2)=\dfrac{1}{17}(-4)^2+\dfrac{2}{17}(-2)^2+\dfrac{3}{17}(0)^2"



"+\dfrac{5}{17}(2)^2+\dfrac{6}{17}(4)^2=\dfrac{140}{17}"


"Var(X)=\sigma^2=E(X^2)-(E(X))^2"


"=\dfrac{140}{17}-(\dfrac{26}{17})^2=\dfrac{1704}{289}"

"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{1704}{289}}=\dfrac{2\sqrt{71}}{17}"